# Interpretable Measurement of CNN Deep Feature Density using Copula and the Generalized Characteristic Function

### v2

@@ 2 Methodology
Our primary goal is to measure, assess, and interpret the statistical distribution of CNN features without making unnecessary and unverified assumptions such as parametric or independence assumptions. The Curse of Dimensionality, as well as the Spatial intuition Limitation complicate this evaluation. In order to measure the distribution under this context, we present a novel approach that combines copula analysis with the Method of Orthogonal Moments. To the best of our knowledge neither of these techniques has been previously applied toward measuring the distribution of CNN features, but they provide a general assumption free way of looking at the high-dimensional interdependence of features and separating this interdepedence from the univariate marginal feature distributions.
@@ 2.1 Copula Analysis
In high-dimensional statistics, copula analysis is a powerful method that allows one to completely separate the marginal distribution of the random variables from their interdependence. Given a set of random variables $\left(X_{1},X_{2},\dots,X_{D}\right)$ and marginal cumulative distribution functions $\left(F_{1},F_{2},\dots,F_{D}\right)$, one can perform a probability integral transform as follows,
@@
The copula $C$ is defined as the cumulative distribution function of the probability integral transform of the random variables as follows,
@@
Moreover the copula density is the probability density function $c\left(y_{1},\;y_{2},\;\dots\;,y_{D}\;\right)$ associated with cumulative copula distribution $C\left(y_{1},\;y_{2},\;\dots\;,y_{D}\;\right)$ as follows,
@@
Typically the probability integral transform converts a marginal distribution into a uniform distribution on the interval $(0,1)$. However, in our approach, we carry out our analysis using a re-scaled version of the probability integral transform to a uniform the interval $(-1,1)$. This is because many well-known orthogonal functions are designed for this interval, and these orthogonal functions allow us to measure the copula density term in greater detail without parametric assumptions. This modified probability integral transform is defined as follows,
$F_{i}(x)=2\;Pr[X_{i}\leq x]-1$ (4)
@@ 2.2 Method of Orthogonal Moments (MOM)
The method of moments are a way of fully describing the shape of a probability distribution through the use of consistent estimators, which asymptotically share sample and population statistics. Assume that $x$ is a finite sample of $n$ elements drawn from infinite population $X$, then a series of well-behaved sample statistics $\hat{\mu}_{t}$ should very closely approximate their population statistics $\mu_{t}$ as follows,
@@
We use the MOM to analyze the copula interdependence $Y_{1},\;\dots\;,Y_{D}$ of deep features separately from the marginals $f(X_{1}),\;\dots\;,f(X_{D})$. As such, we must measure the interdependence between multiple random variables by defining multivariate moments in terms of the expected product of univariate basis statistics in the copula space. If we define $T$ as a $D$ dimensional vector of integers where $T_{d}$ represent the desired moment of the $d^{th}$ random variable $Y_{d}$, then the joint moment $\mu_{T}$ corresponds to inner product of basis vectors as follows,
$\mu_{T}=E\left(\prod_{d=1}^{D}\phi_{T_{d}}(y_{d})\right)\quad\quad\hat{\mu}_{T}=\frac{1}{n}\sum_{i=1}^{n}\left(\prod_{d=1}^{D}\phi_{T_{d}}(y_{d})\right)$ (6)
For our analysis, we specifically define $\phi_{t}$ as either the real-valued Fourier series or the normalized Legendre polynomial series because in addition to being orthogonal over the rescaled copula interval of $\left(-1,1\right)$, these basis functions also have the additional property that all non-constant basis terms exhibit zero integral over $\left(-1,1\right)$, which is useful for our Generalized Characteristic Independence (GCI) metric. The specific forms of these series that we propose also exhibit the property of having unit $L_{2}$ norm over the target interval $(-1,1)$. The Legendre polynomials also have the advantage that, like the power series, $mean$ and $covariance$ are part of the basic shape descriptors, which are highly familiar concepts thereby aiding in practical interpretation.
@@ 2.3 Generalized Characteristic Function
A useful property of the MOM, is that it allows one to recover the actual copula probability density function $c(y)$ in high-dimensions completely non-parametrically, and without any overly-rigid assumptions on the shape of this distribution by means of the Generalized Characteristic Function (GCF). The original Characteristic Function refers to the observation that if one defines the basis set as the Fourier series, then the population moments resembles the Fourier transform of the PDF as follows,
[…]
As such, one can recover the copula density by means of an inverse discrete Fourier transform of the population moments $\mu$. If this process is performed using the sample moments $\hat{\mu}$ then one recovers the sample estimate of the copula density function $\hat{c}$ as follows,
@@
Figure 1: Comparison of orthogonal basis functions over the uniform interval $(-1,1)$. (a) Real-valued Fourier series (b) Normalized Legendre polynomials (c) Legendre polynomials without normalization (d) Chebyshev polynomials.
  • Orthogonal over unit interval $(-1,1)$
@@
  • Unit length $L_{2}$ norm over interval $(-1,1)$
  • All non-constant moments exhibit zero integral over $(-1,1)$
@@ 2.4 Normalized Legendre Polynomials
The Legendre Polynomials figure 1(b,c) are a set of real-valued orthogonal basis functions over the target interval $(-1,1)$ with several desirable properties. Unlike the Chebyshev polynomials figure 1(d), the Legendre polynomials figure 1(c) exhibit zero integral over the interval $(-1,1)$, except trivially for the constant polynomial $P_{0}$. This zero-integral property is highly-desirable for our resultant GCF. The polynomials can be generated efficiently using Bonnet’s recurrence as follows,
@@
$P_{n+1}(y)$ $=\frac{2n+1}{n+1}xP_{n}(y)-\frac{n}{n+1}P_{n-1}(y)$
The Legendre polynomials in this form do not exhibit unit-length $L_{2}$ norm over the interval $(-1,1)$, as such, we propose normalizing the Legendre polynomials based on their $L_{2}$ norm in order to obtain unit-length orthogonal moments as in figure 1(b). This normalized form is obtained as follows,
@@ 2.4.1 Real-valued Fourier Series
As our sample is real-valued, one can equivalently represent the Fourier series as a sum of real-valued $cos$ (even) and $sin$ (odd) harmonic terms. Moreover, it is possible to simplify this to only $cos$ terms if one makes use of the trigonometric phase identity as follows,
$sin(y)=cos(y-\frac{\pi}{2}+2\pi t)\quad\quad\text{for integer }t$ (11)
As such, one can present both even and odd real-valued Fourier basis functions using an elegant and simplified form as follows,
@@
The real-valued Fourier basis functions in this form are shown in figure 1(a). One can see that these basis functions correspond analogously one-to-one with the Legendre and Chebyshev polynomials, with $\phi_{t}$ exhibiting exactly $t$ roots over the target interval $(-1,1)$. This form of the Fourier series also exhibits our desired properties, as the non-constant basis functions have zero integral over the target interval $(-1,1)$. Moreover, for all basis functions the $L_{2}$ norm over the target interval is exactly equal to $1$ when presented in this form.
@@ 2.5 Generalized Characteristic Distance and Independence
Given a set of orthogonal population moments $\mu_{t}$ and $\nu_{t}$ for probability distributions $P$ and $Q$ respectively, the Generalized Characteristic Distance (GCD) fully describes the difference in shape between probability distribution as the Manhattan distance of the Fourier-like transforms of the PDFs. This is calculated by taking the Manhattan distance between the moments as follows,
$D_{char}(P,Q)=$ $\;\sum_{t=1}^{K}\left|\mu_{t}-\nu_{t}\right|\quad$ (13)
In the event that a set of features are completely independent, then the copula density $c\left(y_{1},\;\dots\;y_{D}\right)$ corresponds to the uniform distribution on the hypercube $y_{i}\in(-1,1)$. As such, we new define a Generalized Characteristic Interdependence (GCI) metric $H_{char}(c)$ as the GCD between the copula distribution $c$ and the ideal uniform copula density $Q_{unif}$. $H_{char}$ is exactly zero when variables $y_{1},\;\dots\;,y_{D}$ are statistically independent, and nonzero when these variables show some statistical dependence along one or more orthogonal moments. $H_{char}(c)$ is defined as follows,
@@
The real-valued Fourier and normalized Legendre series have the convenient property that all basis functions integrate to zero over $(-1,1)$ (except trivially the constant basis function $\phi_{0}$). As such, it is straightforward to show that the ideal uniform distribution $Q_{unif}$ has all zero moments. Therefore, for the Fourier and Legendre moments, the GCI simplifies to L1 norm as follows,
$H_{char}(c)=\left|\left|\mu\right|\right|_{1}\quad\quad\quad\text{for Fourier and Legendre moments}$ (15)
@@ 3 Experimental Setup
Figure 2: Illustration of resnet18 (top) and vgg19 (bottom) deep feature layers selected for probability density analysis using copula. Orange shaded regions represent deep feature layers after major architectural blocks that were selected for density analysis. Resnet50 diagram is not shown, but is similar to the resnet18 diagram shown (top).
For our experiments we evaluate three CNN architectures resnet18, resnet50, and vgg19 across four image classification datasets MNIST, cifar10, cifar100, and imagenette2 [5, 16, 4, 12, 10, 30]. Imagenette2 is a subset of Imagenet exhibiting 10 classes with the full resolution images that are center-cropped to $224\times 224$ pixels.
Figure 2 shows the extracted feature spaces from the resnet and vgg architectures. For each image, after every major convolutional block, we obtain a tensor of size $[N_{img},\text{filters},\text{rows},\text{cols}]$. Due to the self-similarity of features in each row and column, we evaluate the distribution for each of the filters over the entire training or test set. Thus we have a sample of $N=N_{img}\times\text{rows}\times\text{cols}$ for every filter for each block.
@@
For imagenette2, we used the pre-trained versions of resnet18, resnet50, and vgg19 as included with PyTorch, with a custom trained final linear classification layer. For mnist, cifar10, and cifar100 the standard versions of resnet and vgg are not designed to work with such small resolution images, and thus we used the small-image optimized versions of these architectures by Kuang [17]. This small-image optimized version is widely used, often without attribution, in papers that achieve high accuracy on these datasets. The small-image resnet and vgg architectures were trained from scratch using a learning rate of 0.01, momentum 0.09, and the SGD optimizer.

### v3

@@ 2.1 Copula Analysis
In high-dimensional statistics, copula analysis is a powerful method that allows one to completely separate the marginal distribution of the random variables from their interdependence. Given a set of random variables $\left(X_{1},X_{2},\dots,X_{D}\right)$ and marginal cumulative distribution functions $\left(F_{1},F_{2},\dots,F_{D}\right)$, one can perform the following probability integral transform.
@@
The copula $C$ is defined as the cumulative distribution function of the probability integral transform of the random variables, as given by the following.
@@
Moreover the copula density is the probability density function $c\left(y_{1},\;y_{2},\;\dots\;,y_{D}\;\right)$ associated with cumulative copula distribution $C\left(y_{1},\;y_{2},\;\dots\;,y_{D}\;\right)$, which is defined by the following.
@@
Typically the probability integral transform converts a marginal distribution into a uniform distribution on the interval $(0,1)$. However, in our approach, we carry out our analysis using a re-scaled version of the probability integral transform to a uniform the interval (1,1)(-1,1). This is because many well-known orthogonal functions are designed for this interval, and these orthogonal functions allow us to measure the copula density term in greater detail without parametric assumptions. This modified probability integral transform is defined by the following.
$F_{d}(x)=2\;Pr[X_{d}\leq x]-1$ (4)
@@ 2.2 Method of Orthogonal Moments (MOM)
The method of moments are a way of fully describing the shape of a probability distribution through the use of consistent estimators, which asymptotically share sample and population statistics. Assume that $x$ is a finite sample of $N$ elements drawn from an infinite population, then the following shows a series of well-behaved sample statistics $\hat{\mu}_{t}$ should very closely approximate their population statistics $\mu_{t}$.
@@
Figure 1: Comparison of orthogonal basis functions over the uniform interval (1,1)(-1,1). (a) Real-valued Fourier series (b) Normalized Legendre polynomials (c) Legendre polynomials without normalization (d) Chebyshev polynomials.
We use the MOM to analyze the copula interdependence $Y_{1},\;\dots\;,Y_{D}$ of deep features separately from the marginals $f(X_{1}),\;\dots\;,f(X_{D})$. As such, we must estimate the interdependence between multiple random variables by defining multivariate moments in terms of the expected product of univariate basis statistics in the copula space. If we define $T$ as a $D$ dimensional vector of integers where $T_{d}$ represent the desired moment of the $d^{th}$ random variable $Y_{d}$, then the joint moment $\mu_{T}$ corresponds to inner product of basis vectors given by the following.
$\mu_{T}=E\left(\prod_{d=1}^{D}\phi_{T_{d}}(Y_{d})\right)\quad\quad\hat{\mu}_{T}=\frac{1}{N}\sum_{i=1}^{N}\left(\prod_{d=1}^{D}\phi_{T_{d}}(Y_{d,i})\right)$ (6)
For our analysis, we specifically define $\phi_{t}$ as either the real-valued Fourier series or the normalized Legendre polynomial series because in addition to being orthogonal over the rescaled copula interval of $\left(-1,1\right)$, these basis functions also have the useful property that all non-constant basis terms exhibit zero integral over $\left(-1,1\right)$, which simplifies our Generalized Characteristic Independence (GCI) metric. The specific forms of these series that we propose also exhibit the property of having unit $L_{2}$ norm over the target interval (1,1)(-1,1). The Legendre polynomials also have the advantage that, like the power series, meanmean and covariancecovariance are part of the basic shape descriptors, which are highly familiar concepts thereby aiding in practical interpretation.
@@ 2.3 Generalized Characteristic Function
A useful property of the MOM, is that it allows one to recover the actual copula probability density function $c(y)$ completely non-parametrically, and without any overly-rigid assumptions on the shape of this distribution by means of the Generalized Characteristic Function (GCF). The original Characteristic Function refers to the observation that if one defines the basis set as the Fourier series, then the following shows how the population moments resembles the Fourier transform of the PDF.
$\mu_{p}=E\left(\phi_{p}(y)\right)=\int_{-\infty}^{\infty}c(y)\;e^{i\pi ty}dy$ (7)
As such, one can recover the copula density by means of an inverse discrete Fourier transform of the population moments $\mu$. If this process is performed using the sample moments $\hat{\mu}$ then using the following, one recovers the sample estimate of the copula density function $\hat{c}$. In the univariate case, this process of recovering the copula density from the characteristic function can be written as follows.
@@
  • Orthogonal over unit interval (1,1)(-1,1)
@@
  • Unit length $L_{2}$ norm over interval (1,1)(-1,1)
  • All non-constant moments exhibit zero integral over (1,1)(-1,1)
@@ 2.4 Normalized Legendre Polynomials
The Legendre Polynomials (figure 1b,c) are a set of real-valued orthogonal basis functions over the target interval (1,1)(-1,1) with several desirable properties. Unlike the Chebyshev polynomials (figure 1d), the Legendre polynomials (figure 1c) exhibit zero integral over the interval (1,1)(-1,1), except trivially for the constant polynomial $P_{0}$. This zero-integral property is highly-desirable for our resultant GCF. The following polynomials can be generated efficiently using Bonnet’s recurrence.
@@
$P_{n+1}(y)$ $=\frac{2n+1}{n+1}yP_{n}(y)-\frac{n}{n+1}P_{n-1}(y)$
The Legendre polynomials in this form do not exhibit unit-length $L_{2}$ norm over the interval (1,1)(-1,1), as such, we propose normalizing the Legendre polynomials based on their $L_{2}$ norm in order to obtain unit-length orthogonal moments as in (figure 1b). This normalized form is obtained from the following.
@@ 2.4.1 Real-valued Fourier Series
As our sample is real-valued, one can equivalently represent the Fourier series as a sum of real-valued coscos (even) and sinsin (odd) harmonic terms. Moreover, it is possible to simplify this to only coscos terms if one makes use of the trigonometric phase identity, which is given by the following.
[…]
As such, one can present both even and odd real-valued Fourier basis functions using an elegant and simplified form, demonstrated by the following.
@@
The real-valued Fourier basis functions in this form are shown in (figure 1a). One can see that these basis functions correspond analogously one-to-one with the Legendre and Chebyshev polynomials, with $\phi_{t}$ exhibiting exactly $t$ roots over the target interval (1,1)(-1,1). This form of the Fourier series also exhibits our desired properties, as the non-constant basis functions have zero integral over the target interval (1,1)(-1,1). Moreover, for all basis functions the $L_{2}$ norm over the target interval is exactly equal to $1$ when presented in this form.
2.5 Multivariate Empirical GCF Estimate of Copula Density
[…]
$\Phi_{T}(\vec{y})=\prod_{d=1}^{D}\phi_{T_{d}}(\vec{y}_{d})$ (13)
As such, the multivariate empirical copula density can be obtained by the inner product of multivariate moments $\hat{\mu}$ and the corresponding orthogonal functions $\Phi$.
$\hat{c}(\vec{y})\quad=\quad\hat{\mu}\;\cdot\;\Phi(\vec{y})\quad=\sum_{T\;\in\;\mathbb{Z}_{K}^{D}}\hat{\mu}_{T}\;\Phi_{T}(\vec{y})\quad$ (14)
@@ 2.6 Generalized Characteristic Distance and Independence
Given a set of orthogonal population moments $\mu_{T}$ and $\nu_{T}$ for probability distributions $P$ and $Q$ respectively, the Generalized Characteristic Distance (GCD) fully describes the difference in shape between probability distribution as the Manhattan distance of the Fourier-like transforms of the PDFs. The following shows how this is calculated by taking the Manhattan distance between the moments.
$D_{char}(P,Q)=$ $\;\sum_{T\;\in\;\mathbb{Z}_{K}^{D}}\left|\mu_{T}-\nu_{T}\right|\quad$ (15)
In the event that a set of features are completely independent, then the copula density $c\left(y_{1},\;\dots\;y_{D}\right)$ corresponds to the uniform distribution on the hypercube $y_{i}\in(-1,1)$. As such, we new define a Generalized Characteristic Interdependence (GCI) metric $H_{char}(c)$ as the GCD between the copula distribution $c$ and the ideal uniform copula density $Q_{unif}$. $H_{char}$ is exactly zero when variables $y_{1},\;\dots\;,y_{D}$ are statistically independent, and nonzero when these variables show some statistical dependence along one or more orthogonal moments. $H_{char}(c)$ is defined by the following.
@@
The real-valued Fourier and normalized Legendre series have the convenient property that all basis functions integrate to zero over (1,1)(-1,1) (except trivially the constant basis function $\phi_{0}$). As such, it is straightforward to show that the ideal uniform distribution $Q_{unif}$ has all zero moments. Therefore, for the Fourier and Legendre moments, the GCI simplifies to L1 norm, as shown in the following.
[…]
2.7 Algorithmic Design
We now describe our algorithm for calculation of a multivariate copula density of deep features based on our methodology. Algorithms 1 and 2 show the calculation of marginal and interdependence terms. Algorithm 3 further shows the estimation of the copula density. Both algorithms depend on additional pseudocode for evaluation of orthogonal functions, as well as a multivariate counter as provided in appendix E.
Algorithm 1 Calculation of copula marginal terms
1:procedure CalculateMargianls(X,D,N)
2:  for $d\leftarrow 1\dots D$ do
3:    $\triangleright\textit{ Calculate Probability Integral Transform}$
4:   for $i\leftarrow 1\dots N$ do
5:     $Y_{d,i}\leftarrow 2Pr[X_{d}\leq X_{d,i}]-1$
6:   end for
7:  end for
8:  $\textbf{return }\left(X,Y\right)$
9:end procedure
Algorithm 2 Calculation of copula interdependence terms
1:procedure CalculateInterdependence(Y,D,N,K)
2:   $\triangleright\textit{ Initialize empirical GCF Density}$
3:  $T=\textsc{Initialize}()$
4:  while $T\neq\text{overflow}$ do
5:   $\hat{\mu}_{T}\leftarrow 0$
6:   $T\leftarrow\textsc{Increment}(T)$
7:  end while
8:
9:   $\triangleright\textit{ Calculate empirical Multivariate Orthogonal Moments}$
10:  for $i\leftarrow 1\dots N$ do
11:
12:    $\triangleright\textit{ Univariate orthogonal functions evaluated at sample }Y_{i}$
13:   $\phi\leftarrow\textsc{EvalOrthogonal}(\ \!Y_{1:D,\ \!i}\ \!)$
14:
15:    $\triangleright\textit{ Multivariate orthogonal moments }\hat{\mu}$
16:   $T=\textsc{Initialize}()$
17:   while $T\neq\text{overflow}$ do
18:     if $\textsc{OrderK}(T)$ then
19:      $\hat{\mu}_{T}\leftarrow\hat{\mu}_{T}+\frac{1}{N}\prod_{d=1}^{D}\phi_{d,T_{d}}$
20:     end if
21:     $T\leftarrow\textsc{Increment}(T)$
22:   end while
23:
24:  end for
25:  $\textbf{return }\hat{\mu}$
26:end procedure
Algorithm 3 Calculation of copula density for $y$
1:procedure CopulaDensity($\vec{y},\hat{\mu}$)
2:  $\hat{c}\leftarrow 0$
3:
4:   $\triangleright\textit{ Evaluate univariate functions at test sample }y$
5:  $\phi\leftarrow\textsc{EvalOrthogonal}(\vec{y})$
6:
7:   $\triangleright\textit{ Multivariate copula density }\hat{c}$
8:  $T=\textsc{Initialize}()$
9:  while $T\neq\text{overflow}$ do
10:   $\Phi_{T}\leftarrow\prod_{d=1}^{D}\phi_{d,T_{d}}$
11:   $\hat{c}\leftarrow\hat{c}+\hat{\mu}_{T}\;\Phi_{T}$
12:   $T\leftarrow\textsc{Increment}(T)$
13:  end while
14:  $\textbf{return }\hat{c}$
15:end procedure
2.8 Formal Definition
Define $X\in\mathbb{R}^{D\times N}$ as the raw sample under consideration for copula analysis. In our case, $X\in\mathbb{R}^{D\times N}$ represents the activation of different $D$ features over a sample of $N$ instances from the training dataset. As such $X_{d}\in\mathbb{R}^{N}$ represents the training sample for the $d^{th}$ feature under consideration.
[…]
Define $\mathcal{G}_{d}$ as the marginal distribution for sample $X_{d}$.
$X_{d}=\left(X_{d,1},X_{d,2},\dots,X_{d,N}\right)\text{ is drawn from marginal }\mathcal{G}_{d}$ (19)
Define $F_{d}(x):\mathbb{R}\rightarrow\mathbb{R}$ as a probability integral transform that rescales values from the marginal distribution of sample $X_{d}$ into a uniform distribution over the interval (1,1)(-1,1).
$F_{d}(x)=2\;Pr[X_{d}\leq x]-1$ (20)
Define $Y\in\mathbb{R}^{D\times N}$ as the sample as transformed by the probability integral transform. As such, $\forall d\in[1,D]\quad\forall i\in[1,N]$,
$Y_{d,i}=F_{d}\left(X_{d,i}\right).$ (21)
Therefore, $\forall d\in[1,D]$, $Y_{d}$ follows a uniform distribution over the target interval (1,1)(-1,1).
$Y_{d}=\left(Y_{d,1},Y_{d,2},\dots,Y_{d,N}\right)\text{ is drawn from }\textit{Uniform}\left(-1,1\right)$ (22)
Define $\phi_{t}(y):\mathbb{R}\rightarrow\mathbb{R}$ as the $t^{th}$ orthogonal function as previously defined in either equation 12 for the Real Fourier series or 10 for the Normalized Legendre series.
[…]
Notice that we consider only the first $K$ functions of the orthogonal series. As such, $t$ belongs to the rank $K$ modular integers, i.e. $t\in\mathbb{Z}_{K}$ because $t$ can only take the values $\{0,1,...,K-1\}$.
[…]
Define $T\in\mathbb{Z}_{K}^{D}$ as a multivariate counter of rank $D$ and modulo $K$. Thus, $T$ allows us to uniquely index each of the multivariate orthogonal functions and moments.
[…]
Define $\Phi_{T}(\vec{y}):\mathbb{R}^{D}\rightarrow\mathbb{R}$ the multivariate orthogonal function of index $T$, which is the product of the orthogonal univariate orthogonal functions.
$\Phi_{T}(\vec{y})=\prod_{d=1}^{D}\phi_{T_{d}}(\vec{y}_{d})$ (26)
Define $\mu_{T}$ as an unobserved multivariate population moment, and $\hat{\mu}_{T}$ as the corresponding multivariate sample moment.
$\mu_{T}=\mathbb{E}\left(\Phi_{T}(Y)\right)\quad\quad\quad\hat{\mu}_{T}=\frac{1}{N}\sum_{i=1}^{N}\left[\;\prod_{d=1}^{D}\phi_{T_{d}}(Y_{d,i})\;\right]$ (27)
Define $\hat{c}(\vec{y}):\mathbb{R}^{D}\rightarrow\mathbb{R}$ as the empirical copula density as evaluated over transformed testing feature vector $\vec{y}$. $c(\vec{y})$ can be calculated by performing the inverse discrete transform from the orthogonal series. Given that our orthogonal series requires the properties of orthogonality over the unit interval (1,1)(-1,1), as well as unit length $L_{2}$ norm over this interval, the inverse is simplifies to the transpose. Therefore the empirical copula density $\hat{c}(\vec{y})$ is recovered by taking the inner product of the set of multivariate moments $\hat{mu}$ with the set of orthogonal functions $\Phi$ evaluated at vector $\vec{y}$.
$\hat{c}(\vec{y})\quad=\quad\hat{\mu}\;\cdot\;\Phi(\vec{y})\quad=\sum_{T\;\in\;\mathbb{Z}_{K}^{D}}\hat{\mu}_{T}\;\Phi_{T}(\vec{y})\quad$ (28)
@@ 3 Experimental Setup
For our experiments we evaluate three CNN architectures ResNet-18, ResNet-50 , and VGG-19 across four image classification datasets MNIST, CIFAR-10, CIFAR-100, and Imagenette2 (deng2012mnist, ; krizhevsky2009learning, ; deng2009imagenet, ; Howard_Imagenette_2019, ; he2016deep, ; simonyan2014very, ). Imagenette2 is a subset of ImageNet exhibiting 10 classes with the full resolution images that are center-cropped to $224\times 224$ pixels.
Figure 2 shows the extracted feature spaces from the ResNet and VGG architectures. For each image, after every major convolutional block, we obtain a tensor of size $N_{img},\text{filters},\text{rows},\text{cols}]$. Due to the self-similarity of features in each row and column, we evaluate the distribution for each of the filters over the entire training or test set. Thus we have a sample of $N=N_{img}\times\text{rows}\times\text{cols}$ for every filter for each block.
@@
For Imagenette2, we used the pre-trained versions of ResNet-18, ResNet-50, and VGG-19 as included with PyTorch, with a custom trained final linear classification layer. For MNIST, CIFAR-10, and CIFAR-100 the standard versions of ResNet and VGG are not designed to work with such small resolution images, and thus we used the small-image optimized versions of these architectures by Kuang kuang2017cifar . This small-image optimized version is widely used, often without attribution, in papers that achieve high accuracy on these datasets. The small-image ResNet and VGG architectures were trained from scratch using a learning rate of 0.01, momentum 0.09, and the SGD optimizer.
@@
Figure 2: Illustration of ResNet-18 (left) and VGG-19 (right) deep feature layers selected for probability density analysis using copula. Orange shaded regions represent deep feature layers after major architectural blocks that were selected for density analysis. ResNet-50 is not shown, but similar to ResNet-18 (left).